OpenStudy (anonymous):

If H(x)= 4x^2-16 was shifted 7 units to the right and 3 down, what would the new equation be? a. h(x)= 4(x+9)^2-17 b. h(x)= 4(x-9)^2-17 c. h(x)= 4(x-7)^2- 19 d. h(x)= 4(x+7)^2 +19

myininaya (myininaya):

i think it is c

myininaya (myininaya):

let me graph it and see what happens

OpenStudy (anonymous):

If the function is shifted to the right then the x becomes x-7. If it is shifted down 3 units then you would subtract 3 H(x) = 4(x-7)^2 -16-3 = 4(x-7)^2 - 19

myininaya (myininaya):

yep it is definitely c